While standing in line for the water fountain, Billy sees his lab partner 3 feet ahead of him and his best friend 9 feet to his right. Billy wants to go ask his lab partner a question, then go chat with his friend, and finally return to the water fountain line. How far will Billy have to walk in all? If necessary, round to the nearest tenth.


1
Expert's answer
2021-12-26T17:18:04-0500

Solution:


Let B be the position of Billy, L be his lab partner's and F be his friend's.

This forms a right-angled triangle.

Apply pythagoras theorem in it.

"BL^2+BF^2=LF^2
\\ \Rightarrow 3^2+9^2=LF^2
\\\Rightarrow 9+81=LF^2
\\ \Rightarrow LF^2=90
\\ \Rightarrow LF=\sqrt{90}\approx9.5\ ft"

Total distance billy has to cover "=BL+BF+LF=3+9+9.5=21.5\ ft"